# A systematic review of metaheuristics-based and machine learning-driven intrusion detection systems in IoT

### v3

@@ 6 Discussions
Multimodal data processing has gained significant momentum with the advent of alignment models, particularly vision-language models, which excel at creating efficient vector-based data structures for storage, retrieval, and utilization. This is especially beneficial in IoT applications where resources such as processing power and memory are limited. Real-time detection leveraging edge and fog computing has also advanced, providing scalable solutions while minimizing latency. Multimodal approaches are crucial for managing diverse datasets and reducing false positives [li2023gasto, xu2024fast, li2024novel]. Emerging research directions include privacy-preserving federated learning, which enables decentralized model training while safeguarding data privacy, and explainable AI, which enhances trust by making IDS outputs interpretable. Additionally, quantum-inspired optimization is being explored for faster convergence, and multi-objective optimization aims to balance detection accuracy, latency, and energy efficiency in IoT environments. In addition, metaheuristic algorithms can provide an alternative to the neural architecture search algorithms for constructing hybrid novel deep learning models, handling multi-objective optimization problems, etc.
7 Conclusion
@@
| [2_rm2020effective] | hybrid PCA-GWO | FS | - | DNN | Kaggle dataset | 99.9 | 95.4 | 100 | ||||||
| [4_khafaga2023voting] | new WOA, guided by DTO | PT | N/A | KNN, RF, NN | RPL-NIDS17 | 95.1 | AUC=99.0, MSE=2.50E-08 | |||||||
| [5_saif2022hiids] | PSO, GA, and DE | FS | 8-10 | KNN,DT | NSL-KDD | 95.71 | ||||||||
| [6_li2022improving] | BOA | FS | - | ANN | NSL-KDD | 93.27 | 94.37 | 92.68 | ||||||
| [7_sanju2023enhancing] | HHO-FDM | FS | - | LSTM +GRU = RNNs | IoT-23 | 98.12 | 98.06 | 98.31 | 98.18 | AUC-ROC=99.82 | ||||
@@
| [8_savanovic2023intrusion] | modified FOA | HPT | N/A | KNN, XGBoost | IoT-healthcare-security-dataset | 99.6997 | 99.6998 | 99.6997 | 99.6996 | |||||
| [9_vaiyapuri2023metaheuristics] | BSA, SCOA | FS, PT | - | FL, KELM | - | 99.45 | 80.26 | 82.67 | 80.95 | |||||
| [11_malibari2022novel] | IAOA, QPSO | FS+ PT | - | DWNN | CICIDS2017 | 98.21 | 96.53 | 98.22 | 97.92 | |||||
| [12_dey2023hybrid] | NSGA-II | FS | 13 | SVM | TON-IoT | 99.48 | ||||||||
@@
| [13_krishna2021attack] | LOA-FOA | FS | - | RF | NBaIoT | 99.86 | 99.94 | 99.94 | 99.86 | FN=7, FP=2 | ||||
| [14_dey2023metaheuristic] | BGSA and BGWO | FS | 4 | DT and EL (AdaBoost and RF) | UNSW-NB15 | 99.41 | 99.92 | 99.33 | 99.09 | 0.03 | ||||
@@
| [17_elmasry2020evolving] | double PSO | FS+HPT | 10 | DNN, LSTM-RNN, and DBN | CICIDS2017 | 95.81 | 95.82 | 95.81 | 95.81 | |||||
@@
| [18_fatani2023enhancing] | MGO, using WOA | FS | - | CNNs | CICIDS-2017 | 99.22 | 99.188 | 99.248 | 99.218 | G-Mean=99.218 | ||||
@@
| [20_alghamdi2022hybrid] | PO | PT | N/A | CFNN | CIDCC-2017 | 99.38 | 99.69 | 99.66 | 99.69 | |||||
| [21_alweshah2022intrusion] | SSO-SA1 and SSO-SA2 | FS | 39.11 | KNN | N-BaIoT | 0.987 | 0.997 | 0.996 | 0.998 | |||||
| [22_anusha2022intrusion] | GSO | FS | - | PCA | NSL-KDD | 93.35 | 91.9 | 94.02 | 95.12 | 2.97 | ||||
| [23_alweshah2023intrusion] | EPC | FS | 38.27 | KNN | N-BaIoT | 98.2 | 99.7 | 99.2 | 99.4 | |||||
| [24_rahmani2024improvement] | GAO-AOA | PT+HPT | N/A | RdNN | TON_IOT | 99.56 | 99.37 | 4 | ||||||
@@
| [25_ghasemi2024new] | GWO | FS+HPT | - | SVM | TON_IOT | 81 | 82 | 84 | 83.57 | |||||
| [26_sagu2023design] | SAEHO, SU-CMO | PT | N/A | CNN+DBN and Bi-LSTM+GRU | UNSW-NB15 | 92.8 | 81 | MCC=0.786, Rand Index = 0.998 (dataset-2) | ||||||
| [27_alkanhel2023hybrid] | GWO-DTO | FS | - | KNN | RPL-NIDS17 | 98.1 | 97.8 | 99.4 | 98.6 | 97.8 | ||||
| [28_forestiero2021metaheuristic] | ABF | HPT | N/A | IoT2Vec | CASAS dataset | 92.98 | avg entropy= 0.7478 | |||||||
| [29_bahaa2022novel] | APSO-WOA | HPT | N/A | CNN | N-BaIoT | 94.54 | 95.2 | kappa= 0.936, hamming loss= 0.944 , JSC= 0.9 | ||||||
| [30_davahli2020hybridizing] | GA-GWO | FS | 92 | SVM | AWID | 99.1 | 96.03 | 97.64 | 99.32 | 0.69 | ||||
| [31_kumar2022intellectual] | HHGS-ROA | FS | - | SVM | AWID | 99.16 | 99.76 | 99.4 | 99.58 | 0.2 | MCC=99.97 | |||
| [32_habib2020modified] | MOPSO-Lévy | FS | 44.33 | KNN | N-BaIoT | 97.06 | 88.69 | 1.66 | TPR=0.7506, TNR=0.9834, G-mean=0.8317, AUC =0.867 | |||||
@@
| [34_stankovic2022feature] | a hybrid ABC | FS | 14.8, 11.9 | ELM | UNSW-NB15 | 71.54 | 80.58 | |||||||
| [35_baniasadi2022novel] | NSBPSO | PT | N/A | DCNN | UNSW-NB15 and Bot-IoT | 98.86 | 99.03 | 95.32 | MSE=0.00053 | |||||
| [36_aljebreen2023binary] | BCOA, SCA | FS, PT+HPT | - | CCR-ELM | WSN-DS | 99.63 | 97.91 | 94.52 | 99.67 | |||||
| [37_lv2023binary] | BBFA | FS | 23 | SVM | N-BaIoT | 99.2 | 99 | 0.006 | ||||||
| [38_gaber2023industrial] | PSO-BA | FS | 16 | RF | WUSTL-IIOT-2021 | 99.99 | 99.6 | 99.6 | 99.6 | |||||
| [39_vanitha2023improved] | IACO | FS | - | EL using DDT, ANFIS and MDSVM | UNSW-NB15 | 97.375 | 92.365 | 6.67 | ||||||
| [40_alruwaili2023red] | RKOA, LCWOA | FS,HPT | - | EL using LSTM, BiLSTM, and BiGRU | WSN-DS | 98.94 | 75.33 | 79.52 | 75.33 | AUC=85.48 | ||||
@@
| [41_ethala2022hybrid] | SMO-Hierarchical PSO(HPSO) | FS | 22 | RF | NSL-KDD | B | 98.31 | 98.61 | 98.41 | 98.56 | 0.21 | AUC=99.87, MCC=0.17 | ||
@@
| [42_asgharzadeh2023anomaly] | BMECapSA | FS | 12,18 | CNN | TON-IoT | 99.99 | 99.99 | 99.99 | 99.99 | 0.0001 | FNR=0.00002 | |||
| [43_jayasankar2024intrusion] | DFWA | FS | - | ODRNN | IDS dataset | 96.11 | 96.11 | 97.21 | 97.31 | 3.03 | ||||
| [44_keserwani2021smart] | GWO-PSO | FS | - | RF | KDDCup-99, NSL–KDD, CICIDS-2017 | M | 99.88, 99.25, 99.87 | |||||||
| [45_karthikeyan2024firefly] | GWO, FOA | PT, FS | - | SVM | NSL-KDD | 99.29 | 98.12 | 96.23 | 99.59 | AUC=98.51 | ||||
@@
| [46_jayalatchumy2024improved] | enhanced CrSA | FS | 11,7 | EL | UNSW-NB15 | 97.75 | 83.57 | 83.39 | 81.66 | |||||
@@
| [47_chander2023metaheuristic] | new DHOAF, SpSO | FS,PT | - | CRNN | UCI SECOM | 97.88 | 92.42 | 89.87 | 91.1 | |||||
| [48_sarwar2022design] | PSO | FS | 17 | RF | IoTID20 | 98 (B), | ||||||||
@@
| [49_fatani2021advanced] | AQUO | FS | 10 | CNN | KDDCup-99 | B | 99.922 | 94.283 | 92.256 | 92.683 | ||||
@@
| [50_jovanovic2022feature] | SCSO | FS | 14.5,11.7 | ELM | CICIDS-2017 | 98.7 | 96.17 | |||||||
@@
| [52_pan2021lightweight] | Compact SCA | HPT | N/A | kNN | UNSW-NB15 | 98.27 | 97.94 | 5.82 | ||||||
| [53_qaddoura2024evolving] | GWO, PSO, MVO | PT+HPT | N/A | RWNN | IoTID20 | G-mean=0.7283 | ||||||||
| [54_jeyaselvi2023highly] | EXPSO-STFA | FS | - | LAANN | KDDCup-99, NSL-KDD, | 95.65 | 94.74 | 93.54 | 95.64 | 92.74 | 14.52 | FNR=10.2 | ||
@@
| [56_xu2023iot] | BGWO | FS | - | XGBoost | WUSTL-EHMS-2020 | M | 98.897 | 98.8923 | 98.897 | 98.8846 | ||||
@@
| [57_prajisha2022efficient] | ECSSA | FS | - | LightGBM | MQTTset | 98.35 | 97.38 | 97.68 | 98.56 | |||||
| [58_nazir2023novel] | TS | FS | 13 | EL using RF | TON_IoT | 99.5 | 97.92 | 0.004 | ||||||
| [59_alamiedy2020anomaly] | multi-objective GWO | FS | 4 | SVM | NSL–KDD | 87.59 | ||||||||
@@
| [60_gangula2023intrusion] | FOA | FS | - | EL using KNN,SVM,LSTM,MLP | NSL-KDD | B | 98.41 | 98.68 | 98.46 | 98.68 | 0.24 | AUC=99.79, MCC=0.26 | ||
| [61_davahli2020lightweight] | GABGWO | FS | 94 | SVM | AWID and KDDCup-99 | 99.09 | 96.31 | 99.3 | 97.84 | 0.68 | ||||
@@
| [65_anushiya2023new] | AAFSO,GA | FS,HPT | - | FR-CNN | BoT-IoT | 93.7756 | 86.6687 | 95.874 | 91.0393 | |||||
| [66_gadekallu2023moth] | MFO | FS | 14 | EL using LR, RF and XGBoost | UNSW-NB15 | 100 | 99.5 | 100 | ||||||
| [67_gupta2022hybrid] | HCSGA | FS | - | DLHNN | NSL-KDD | 99.52 | 97.55 | 97.16 | 96.78 | |||||
| [68_sandhya2021enhancing] | SMO | FS | - | RF | NSL-KDD | 99.675 | 99.955 | 99.9425 | 99.9325 | AUC=99.3025 | ||||
| [69_kannan2020intrusion] | ACO | FS | - | PCA | KDDCup-99 | 91 | 1.8 | |||||||
| [70_bathula2022designing] | SHO | HPT | N/A | ANN | KDDcup99 | 98.16 | 98.06 | 98.03 | 98.04 | 98.27 | 1.73 | FNR=1.97, NPV=98.27, MCC=96.30 | ||
@@
| [71_basheri2023quantum] | QCSO, HS | Clustering, HPT | N/A | CRNN | KDDCup-99 | 92.04 | 6.86 | |||||||
| [72_ramana2022wogru] | WOA | HPT | N/A | GRU | WSN-DS | 99.804 | 99.868 | 99.83 | 99.866 | 99.826 | ||||
@@
| [74_saheed2023novel] | HAEMPSO,modified PSO | FS,HPT | - | DNN | UNSW-NB15 | 94.62 | 93.8 | |||||||
@@
| [76_shahapure2021water] | WWO-MSO (WMSA) | FS | - | DRNN | KDDCup-99 | 94.5 | 92.9 | 96.4 | ||||||
| [77_om2021harmony] | HSHO | FS | - | DRL | 96.925 | TPR=96.9, TNR=97.92 | ||||||||
@@
| [78_gopalakrishnan2022new] | AF-EFO | FS+PT | - | HR-OELM using DNN,RF,Adaboost | CICIDS-2018 | 95.3 | 99.7 | 90.8 | 94.9 | 83.6 | 1.5 | FNR=05.7, NPV=83.6, MCC=065.4 | ||
@@
| [79_mohy2023whale] | WOA | FS | - | RBFNN | Merged | 95.93 | MCC=82.68 | |||||||
| [80_bajpai2024hybrid] | WOA | HPT | N/A | XGBoost (EGB) | IoTID20 | 98.86 | 98.67 | 99.91 | 99.30 | AUC=98.91 | ||||
@@
| [81_hanafi2024intrusion] | GJO-OBL (IBGJO) | FS | 32,20 | LSTM | NSL-KDD | 98.93 | 97.98 | 98.37 | 98.17 | |||||
@@
| [82_kethineni2024intrusion] | WHO | PT | N/A | a fused CNN model with Bi-GRU | ToN-IoT | 99.71 | 99.89 | 99.05 | 99.02 | |||||
| [83_singh2024intrusion] | CO-IHHO | FS | - | DT and KNN | BoT-IoT | 100(B), | ||||||||
@@
| [84_mahanipour2024enhancing] | hybrid BQABC-GA | FS | 11,10.6,10.33 | KNN | BoT-IoT | 98.49 | 99.79 | 99.27 | 0.73 | |||||
| [85_vijayan2024original] | TuSO | HPT | N/A | EL combining RF, XGBoost, LightGBM (LGBM), and CatBoost | MQTT dataset | 99.12 | 97.89 | 95.24 | 96.37 | |||||
@@
| [86_aburasain2024enhanced] | Enhanced BWO, BES | HPT,FS | - | HDL | Edge-IIoTset | 98.35 | 84.85 | 80.95 | 82.79 | |||||
| [87_harbi2024bio] | a modified FOA | FS | 39 | DT | Edge-IIoT | 79.64 | ||||||||
@@
| [88_rabie2024novel] | DRFO | FS | - | DBRF | UNSW-NB 15 | 98.5 | 99 | 99 | 98.5 | 8.2 | ||||
| [89_elsedimy2024novel] | Improved GWO (IGWO) | HPT | N/A | QSVM | BoT-IoT | 99.11 | 99.45 | 99.34 | 97.48 | |||||
| [90_biju2024evaluated] | EBSA | PT | N/A | DBN (RBMs,MLPs) | NSL-KDD | 98.96 | 99.4 | 98.87 | ||||||
@@
| [92_sharmamulti] | multi-objective PDO | FS | 10-9,11-15,12 | SVM,KNN | IoTID20 | KNN | 99.402 | 99.386 | 0.006 | |||||
| [93_cherian2024iot] | BOA | PT | N/A | DBN | UNSW-NB15 | 97.77 | 96.62 | 93.85 | 91.77 | MCC=91.23 | ||||
| [94_aljehane2024golden] | a new GJOA, SSA | FS, HPT | - | A-BiLSTM | CICIDS-2017 | 99.69 | 98.92 | 98.92 | 98.92 | MCC=98.74 | ||||
@@
| [96_vadakkethil2024mayfly] | MOA | FS | - | BiLSTM | UNSW-NB15 | 99.35 | 98.49 | 99.28 | 98.64 | |||||
| [97_10.1007/s11276-023-03435-0] | TLBO | FS | - | RF | UNSW-NB15 | 86.5 | ||||||||
@@
| [98_moizuddin2022bio] | Generalized Mean GWO | FS | 15,5 | ElasticNet Contractive AE | BoT-IoT | M | 99.99 | 99.5 | AUC=100, MCC=99.51 | |||||
| [99_kaviarasan2023network] | COA | PT | N/A | 1D-CNN+COA (creating a HNM) | NSL-KDD | 87.19 | 88.28 | 89.49 | 91.19 | |||||
@@
| [101_deore2022hybrid] | CCSO | PT | N/A | Deep LSTM | BoT-IoT | 96.71 | 96.35 | 91.985 | ||||||
| [102_makhadmeh2024intrusion] | modified AOA | FS | - | KNN | BoT-IoT | 99.998 | ||||||||
@@
| [105_10541183] | combination of IGC and SA | FS | - | an AE with DNN | CICIDS2017 | 99.4 | 99.47 | 99.34 | 99.4 | |||||
@@
| [106_escorcia2023sea] | STFA,SpSO | FS,HPT | - | DBN | TON_IoT | 99.51 | 99.73 | |||||||
| [107_10.1016/j.comcom.2023.03.005] | MPO | FS | - | RNN | NSL-KDD | TNR=94, TPR=94 | ||||||||
@@
| [111_dakic2024intrusion] | GSAPSO | HPT | N/A | XGBoost and KNN | CAN dataset | (KNN) | 79.10 | 82.9847 | 79.0996 | 70.8823 | ||||
| [112_latif2024dtl] | GA | HPT | N/A | EL (CNNs-based) | Edge_IIoTset | M | 100 | 100 | 100 | 100 | Cohen’s Kappa score=100 | |||
@@
| [114_alamro2023modeling] | ALO, FPA | FS, HPT | - | CNN+LSTM | ToN-IoT | 99.31 | 99.31 | 99.31 | 99.31 | AUC=99.31 | ||||
| [115_otoum2021ids] | HMS | clustering | N/A | LightNet, Deep Q-learning | NSL-KDD | 96.6 | 96.8 | 96.9 | ||||||

### v4

@@ VI Discussions
Multimodal data processing has gained significant momentum with the advent of alignment models, particularly vision-language models, which excel at creating efficient vector-based data structures for storage, retrieval, and utilization. This is especially beneficial in IoT applications where resources such as processing power and memory are limited. Real-time detection leveraging edge and fog computing has also advanced, providing scalable solutions while minimizing latency. Multimodal approaches are crucial for managing diverse datasets and reducing false positives [110, 174, 107]. Emerging research directions include privacy-preserving federated learning, which enables decentralized model training while safeguarding data privacy, and explainable AI, which enhances trust by making IDS outputs interpretable. Additionally, quantum-inspired optimization is being explored for faster convergence, and multi-objective optimization aims to balance detection accuracy, latency, and energy efficiency in IoT environments. In addition, metaheuristic algorithms can provide an alternative to the neural architecture search algorithms for constructing hybrid novel deep learning models, handling multi-objective optimization problems, etc.
VII Conclusion
@@
| [142] | hybrid PCA-GWO | FS | - | DNN | Kaggle dataset | 99.9 | 95.4 | 100 | ||||||
| [98] | new WOA, guided by DTO | PT | N/A | KNN, RF, NN | RPL-NIDS17 | 95.1 | AUC=99.0, MSE=2.50E-08 | |||||||
| [147] | PSO, GA, and DE | FS | 8-10 | KNN,DT | NSL-KDD | 95.71 | ||||||||
| [109] | BOA | FS | - | ANN | NSL-KDD | 93.27 | 94.37 | 92.68 | ||||||
| [150] | HHO-FDM | FS | - | LSTM +GRU = RNNs | IoT-23 | 98.12 | 98.06 | 98.31 | 98.18 | AUC-ROC=99.82 | ||||
@@
| [152] | modified FOA | HPT | N/A | KNN, XGBoost | IoT-healthcare-security-dataset | 99.6997 | 99.6998 | 99.6997 | 99.6996 | |||||
| [166] | BSA, SCOA | FS, PT | - | FL, KELM | - | 99.45 | 80.26 | 82.67 | 80.95 | |||||
| [116] | IAOA, QPSO | FS+ PT | - | DWNN | CICIDS2017 | 98.21 | 96.53 | 98.22 | 97.92 | |||||
| [46] | NSGA-II | FS | 13 | SVM | TON-IoT | 99.48 | ||||||||
@@
| [103] | LOA-FOA | FS | - | RF | NBaIoT | 99.86 | 99.94 | 99.94 | 99.86 | FN=7, FP=2 | ||||
| [45] | BGSA and BGWO | FS | 4 | DT and EL (AdaBoost and RF) | UNSW-NB15 | 99.41 | 99.92 | 99.33 | 99.09 | 0.03 | ||||
@@
| [54] | double PSO | FS+HPT | 10 | DNN, LSTM-RNN, and DBN | CICIDS2017 | 95.81 | 95.82 | 95.81 | 95.81 | |||||
@@
| [59] | MGO, using WOA | FS | - | CNNs | CICIDS-2017 | 99.22 | 99.188 | 99.248 | 99.218 | G-Mean=99.218 | ||||
@@
| [9] | PO | PT | N/A | CFNN | CIDCC-2017 | 99.38 | 99.69 | 99.66 | 99.69 | |||||
| [17] | SSO-SA1 and SSO-SA2 | FS | 39.11 | KNN | N-BaIoT | 0.987 | 0.997 | 0.996 | 0.998 | |||||
| [21] | GSO | FS | - | PCA | NSL-KDD | 93.35 | 91.9 | 94.02 | 95.12 | 2.97 | ||||
| [18] | EPC | FS | 38.27 | KNN | N-BaIoT | 98.2 | 99.7 | 99.2 | 99.4 | |||||
| [139] | GAO-AOA | PT+HPT | N/A | RdNN | TON_IOT | 99.56 | 99.37 | 4 | ||||||
@@
| [69] | GWO | FS+HPT | - | SVM | TON_IOT | 81 | 82 | 84 | 83.57 | |||||
| [144] | SAEHO, SU-CMO | PT | N/A | CNN+DBN and Bi-LSTM+GRU | UNSW-NB15 | 92.8 | 81 | MCC=0.786, Rand Index = 0.998 (dataset-2) | ||||||
| [14] | GWO-DTO | FS | - | KNN | RPL-NIDS17 | 98.1 | 97.8 | 99.4 | 98.6 | 97.8 | ||||
| [62] | ABF | HPT | N/A | IoT2Vec | CASAS dataset | 92.98 | avg entropy= 0.7478 | |||||||
| [26] | APSO-WOA | HPT | N/A | CNN | N-BaIoT | 94.54 | 95.2 | kappa= 0.936, hamming loss= 0.944 , JSC= 0.9 | ||||||
| [42] | GA-GWO | FS | 92 | SVM | AWID | 99.1 | 96.03 | 97.64 | 99.32 | 0.69 | ||||
| [104] | HHGS-ROA | FS | - | SVM | AWID | 99.16 | 99.76 | 99.4 | 99.58 | 0.2 | MCC=99.97 | |||
| [75] | MOPSO-Lévy | FS | 44.33 | KNN | N-BaIoT | 97.06 | 88.69 | 1.66 | TPR=0.7506, TNR=0.9834, G-mean=0.8317, AUC =0.867 | |||||
@@
| [161] | a hybrid ABC | FS | 14.8, 11.9 | ELM | UNSW-NB15 | 71.54 | 80.58 | |||||||
| [28] | NSBPSO | PT | N/A | DCNN | UNSW-NB15 and Bot-IoT | 98.86 | 99.03 | 95.32 | MSE=0.00053 | |||||
| [11] | BCOA, SCA | FS, PT+HPT | - | CCR-ELM | WSN-DS | 99.63 | 97.91 | 94.52 | 99.67 | |||||
| [112] | BBFA | FS | 23 | SVM | N-BaIoT | 99.2 | 99 | 0.006 | ||||||
| [65] | PSO-BA | FS | 16 | RF | WUSTL-IIOT-2021 | 99.99 | 99.6 | 99.6 | 99.6 | |||||
| [167] | IACO | FS | - | EL using DDT, ANFIS and MDSVM | UNSW-NB15 | 97.375 | 92.365 | 6.67 | ||||||
| [16] | RKOA, LCWOA | FS,HPT | - | EL using LSTM, BiLSTM, and BiGRU | WSN-DS | 98.94 | 75.33 | 79.52 | 75.33 | AUC=85.48 | ||||
@@
| [57] | SMO-Hierarchical PSO(HPSO) | FS | 22 | RF | NSL-KDD | B | 98.31 | 98.61 | 98.41 | 98.56 | 0.21 | AUC=99.87, MCC=0.17 | ||
@@
| [24] | BMECapSA | FS | 12,18 | CNN | TON-IoT | 99.99 | 99.99 | 99.99 | 99.99 | 0.0001 | FNR=0.00002 | |||
| [86] | DFWA | FS | - | ODRNN | IDS dataset | 96.11 | 96.11 | 97.21 | 97.31 | 3.03 | ||||
| [96] | GWO-PSO | FS | - | RF | KDDCup-99, NSL–KDD, CICIDS-2017 | M | 99.88, 99.25, 99.87 | |||||||
| [92] | GWO, FOA | PT, FS | - | SVM | NSL-KDD | 99.29 | 98.12 | 96.23 | 99.59 | AUC=98.51 | ||||
@@
| [85] | enhanced CrSA | FS | 11,7 | EL | UNSW-NB15 | 97.75 | 83.57 | 83.39 | 81.66 | |||||
@@
| [35] | new DHOAF, SpSO | FS,PT | - | CRNN | UCI SECOM | 97.88 | 92.42 | 89.87 | 91.1 | |||||
| [151] | PSO | FS | 17 | RF | IoTID20 | 98 (B), | ||||||||
@@
| [60] | AQUO | FS | 10 | CNN | KDDCup-99 | B | 99.922 | 94.283 | 92.256 | 92.683 | ||||
@@
| [89] | SCSO | FS | 14.5,11.7 | ELM | CICIDS-2017 | 98.7 | 96.17 | |||||||
@@
| [131] | Compact SCA | HPT | N/A | kNN | UNSW-NB15 | 98.27 | 97.94 | 5.82 | ||||||
| [136] | GWO, PSO, MVO | PT+HPT | N/A | RWNN | IoTID20 | G-mean=0.7283 | ||||||||
| [87] | EXPSO-STFA | FS | - | LAANN | KDDCup-99, NSL-KDD, | 95.65 | 94.74 | 93.54 | 95.64 | 92.74 | 14.52 | FNR=10.2 | ||
@@
| [173] | BGWO | FS | - | XGBoost | WUSTL-EHMS-2020 | M | 98.897 | 98.8923 | 98.897 | 98.8846 | ||||
@@
| [134] | ECSSA | FS | - | LightGBM | MQTTset | 98.35 | 97.38 | 97.68 | 98.56 | |||||
| [127] | TS | FS | 13 | EL using RF | TON_IoT | 99.5 | 97.92 | 0.004 | ||||||
| [7] | multi-objective GWO | FS | 4 | SVM | NSL–KDD | 87.59 | ||||||||
@@
| [67] | FOA | FS | - | EL using KNN,SVM,LSTM,MLP | NSL-KDD | B | 98.41 | 98.68 | 98.46 | 98.68 | 0.24 | AUC=99.79, MCC=0.26 | ||
| [41] | GABGWO | FS | 94 | SVM | AWID and KDDCup-99 | 99.09 | 96.31 | 99.3 | 97.84 | 0.68 | ||||
@@
| [23] | AAFSO,GA | FS,HPT | - | FR-CNN | BoT-IoT | 93.7756 | 86.6687 | 95.874 | 91.0393 | |||||
| [66] | MFO | FS | 14 | EL using LR, RF and XGBoost | UNSW-NB15 | 100 | 99.5 | 100 | ||||||
| [74] | HCSGA | FS | - | DLHNN | NSL-KDD | 99.52 | 97.55 | 97.16 | 96.78 | |||||
| [149] | SMO | FS | - | RF | NSL-KDD | 99.675 | 99.955 | 99.9425 | 99.9325 | AUC=99.3025 | ||||
| [90] | ACO | FS | - | PCA | KDDCup-99 | 91 | 1.8 | |||||||
| [30] | SHO | HPT | N/A | ANN | KDDcup99 | 98.16 | 98.06 | 98.03 | 98.04 | 98.27 | 1.73 | FNR=1.97, NPV=98.27, MCC=96.30 | ||
@@
| [29] | QCSO, HS | Clustering, HPT | N/A | CRNN | KDDCup-99 | 92.04 | 6.86 | |||||||
| [140] | WOA | HPT | N/A | GRU | WSN-DS | 99.804 | 99.868 | 99.83 | 99.866 | 99.826 | ||||
@@
| [146] | HAEMPSO,modified PSO | FS,HPT | - | DNN | UNSW-NB15 | 94.62 | 93.8 | |||||||
@@
| [154] | WWO-MSO (WMSA) | FS | - | DRNN | KDDCup-99 | 94.5 | 92.9 | 96.4 | ||||||
| [129] | HSHO | FS | - | DRL | 96.925 | TPR=96.9, TNR=97.92 | ||||||||
@@
| [71] | AF-EFO | FS+PT | - | HR-OELM using DNN,RF,Adaboost | CICIDS-2018 | 95.3 | 99.7 | 90.8 | 94.9 | 83.6 | 1.5 | FNR=05.7, NPV=83.6, MCC=065.4 | ||
@@
| [123] | WOA | FS | - | RBFNN | Merged | 95.93 | MCC=82.68 | |||||||
| [27] | WOA | HPT | N/A | XGBoost (EGB) | IoTID20 | 98.86 | 98.67 | 99.91 | 99.30 | AUC=98.91 | ||||
@@
| [77] | GJO-OBL (IBGJO) | FS | 32,20 | LSTM | NSL-KDD | 98.93 | 97.98 | 98.37 | 98.17 | |||||
@@
| [97] | WHO | PT | N/A | a fused CNN model with Bi-GRU | ToN-IoT | 99.71 | 99.89 | 99.05 | 99.02 | |||||
| [159] | CO-IHHO | FS | - | DT and KNN | BoT-IoT | 100(B), | ||||||||
@@
| [114] | hybrid BQABC-GA | FS | 11,10.6,10.33 | KNN | BoT-IoT | 98.49 | 99.79 | 99.27 | 0.73 | |||||
| [170] | TuSO | HPT | N/A | EL combining RF, XGBoost, LightGBM (LGBM), and CatBoost | MQTT dataset | 99.12 | 97.89 | 95.24 | 96.37 | |||||
@@
| [5] | Enhanced BWO, BES | HPT,FS | - | HDL | Edge-IIoTset | 98.35 | 84.85 | 80.95 | 82.79 | |||||
| [78] | a modified FOA | FS | 39 | DT | Edge-IIoT | 79.64 | ||||||||
@@
| [137] | DRFO | FS | - | DBRF | UNSW-NB 15 | 98.5 | 99 | 99 | 98.5 | 8.2 | ||||
| [55] | Improved GWO (IGWO) | HPT | N/A | QSVM | BoT-IoT | 99.11 | 99.45 | 99.34 | 97.48 | |||||
| [34] | EBSA | PT | N/A | DBN (RBMs,MLPs) | NSL-KDD | 98.96 | 99.4 | 98.87 | ||||||
@@
| [156] | multi-objective PDO | FS | 10-9,11-15,12 | SVM,KNN | IoTID20 | KNN | 99.402 | 99.386 | 0.006 | |||||
| [37] | BOA | PT | N/A | DBN | UNSW-NB15 | 97.77 | 96.62 | 93.85 | 91.77 | MCC=91.23 | ||||
| [12] | a new GJOA, SSA | FS, HPT | - | A-BiLSTM | CICIDS-2017 | 99.69 | 98.92 | 98.92 | 98.92 | MCC=98.74 | ||||
@@
| [165] | MOA | FS | - | BiLSTM | UNSW-NB15 | 99.35 | 98.49 | 99.28 | 98.64 | |||||
| [93] | TLBO | FS | - | RF | UNSW-NB15 | 86.5 | ||||||||
@@
| [124] | Generalized Mean GWO | FS | 15,5 | ElasticNet Contractive AE | BoT-IoT | M | 99.99 | 99.5 | AUC=100, MCC=99.51 | |||||
| [94] | COA | PT | N/A | 1D-CNN+COA (creating a HNM) | NSL-KDD | 87.19 | 88.28 | 89.49 | 91.19 | |||||
@@
| [43] | CCSO | PT | N/A | Deep LSTM | BoT-IoT | 96.71 | 96.35 | 91.985 | ||||||
| [115] | modified AOA | FS | - | KNN | BoT-IoT | 99.998 | ||||||||
@@
| [36] | combination of IGC and SA | FS | - | an AE with DNN | CICIDS2017 | 99.4 | 99.47 | 99.34 | 99.4 | |||||
@@
| [56] | STFA,SpSO | FS,HPT | - | DBN | TON_IoT | 99.51 | 99.73 | |||||||
| [48] | MPO | FS | - | RNN | NSL-KDD | TNR=94, TPR=94 | ||||||||
@@
| [40] | GSAPSO | HPT | N/A | XGBoost and KNN | CAN dataset | (KNN) | 79.10 | 82.9847 | 79.0996 | 70.8823 | ||||
| [106] | GA | HPT | N/A | EL (CNNs-based) | Edge_IIoTset | M | 100 | 100 | 100 | 100 | Cohen’s Kappa score=100 | |||
@@
| [8] | ALO, FPA | FS, HPT | - | CNN+LSTM | ToN-IoT | 99.31 | 99.31 | 99.31 | 99.31 | AUC=99.31 | ||||
| [130] | HMS | clustering | N/A | LightNet, Deep Q-learning | NSL-KDD | 96.6 | 96.8 | 96.9 | ||||||